OpenStudy (anonymous):

Which of the following statements says that a number is between -3 and 3? |x| = 3 |x| < 3 |x| > 3

OpenStudy (anonymous):

the first one

OpenStudy (jdoe0001):

\(\large { |x|=3\implies \begin{cases} +x=3\implies x=3\\ -x=3\implies x=-3 \end{cases}\implies x=\pm 3 \\ \quad \\ |x|<3\implies \begin{cases} +x<3\implies x<3\\ -x<3\implies x>-3 \end{cases}\implies -3<x<3 \\ \quad \\ |x|>3\implies \begin{cases} +x>3\implies x>3\\ -x>3\implies x<-3 \end{cases}\implies 3 < x < -3 } \) recall that when multiplying an inequality by a negative value, you'd need to flip the inequality sign any ideas?